OpenStudy (anonymous):

do you use calculus or precalulus to find the distance traveled in 15 sec by an object traveling at a constant velocity of v(t)=20+7cos t ft/sec

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

but why?

OpenStudy (anonymous):

because v(t)=20+7cos t is not a constant function

OpenStudy (anonymous):

ok but do you use calculus or precalulus? a bicyclist is riding on a path modeled by the equation y=0.04(8x-x^2), where x and y are measured in miles. find the rate of change of evelation when x=2

OpenStudy (anonymous):

y=0.04(8x-x^2) is not a linear equation hence you can't find slope w/o using calculus

OpenStudy (anonymous):

and how about a bicyclist is riding on a path modeled by the equation y=0.08x where x and y are measured in miles.Find the rate of change of the elevation when x=2. Since this is linear I can use pre cal? just y over x?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok thank youuu so much